OpenStudy (anonymous):

Which of the following functions are solutions of the differential equation y'' + y = 5 sin x? (How do I solve this?)

OpenStudy (anonymous):

The list of potential answers are y = − 5/2x cos x y = 5 sin x y = 5/2x sin x y = 5x sin x − 2x cos x y = 5 cos x

OpenStudy (anonymous):

Do I just derive each of the answers twice and then plug it in to the equation and see if it equals?